# Analytic Langlands correspondence from SoV

### v1

@@ 5.1 Definitions
$\Theta^{(n,i)}_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[3.22916pt] x_{1}&\dots&x_{n}\end{smallmatrix}\Big)=\frac{1}{\pi^{n}}\int\prod_{r=1}^{n}\frac{d^{2}k_{r}\,e^{2\mathrm{i}\,\mathrm{Im}(k_{r}x_{r})}}{|k_{r}|^{2j_{r}+2}}\prod_{a=0}^{i}\delta^{(2)}\big({\textstyle\sum_{r=1}^{n}z_{r}^{a}k_{r}}\big)$ (5.45)
$\hskip 184.9429pt\times|\hat{\rho}_{i}(\bm{k})|^{2}\prod_{k=1}^{n-2-i}\chi\big(y_{k}(\bm{k}),\bar{y}_{k}(\bm{k})\big),$
with $y_{1}(\bm{k}),\dots,y_{n-2-i}(\bm{k})$ and $\hat{\rho}_{i}(\bm{k})$, $\bm{k}=(k_{1},\dots,k_{n})$, defined by
$k_{r}=\hat{\rho}_{i}(\bm{k})\frac{\prod_{k=1}^{n-2-i}(z_{r}-y_{k}(\bm{k}))}{\prod_{s\neq r}(z_{r}-z_{s})},\qquad r=1,\dots,n,\qquad\hat{\rho}_{i}(\bm{k})=\sum_{r=1}^{n}z_{r}^{1+i}k_{r},$ (5.46)
for $i=0,1$, respectively. The main differences between $i=0$ and $i=1$ are a larger range of values for the indices $y_{k}$ in the case $i=0$, and the extra delta-distribution $\delta^{(2)}\big({\textstyle\sum_{r=1}^{n}z_{r}k_{r}}\big)$, in the case $i=1$.
The functions $\Theta^{(n,i)}_{\chi}$ will turn out to be proportional to the functions $\Sigma^{(n,i)}_{\chi}\big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[1.07639pt] x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\big)$ defined by $\big(\mathsf{S}_{\bm{j},i}^{(n)}\big)^{-1}\phi_{\chi}$ in the limit $z_{n}\to\infty$.
The integrals on the right side of (LABEL:Fourier(n)) are initially defined in the distributional sense. In order to see that (LABEL:Fourier(n)) actually defines smooth functions away from singular divisors of co-dimension one, we may consider the inverse $\mathsf{SoV}$-transformation in the case where the number $n$ of punctures is replaced by $m=n+1$, with punctures $z_{r}$ now labelled by an index $r$ running from $1$ to $n+1$, assuming $z_{n+1}=\infty$. The inverse $\mathsf{SoV}$-transformation will in this case act on products of functions $\hat{\chi}$ representing real opers on the $(n+1)$-punctured sphere. As noted above, it will define a function denoted as $\Sigma^{(n+1)}_{\hat{\chi}}\big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[2.15277pt] x_{1}&\dots&x_{n}\end{smallmatrix};j_{n+1}\big)$ away from singular divisors of co-dimension one. We furthermore know that $\Sigma^{(n+1)}_{\hat{\chi}}$ is meromorphic in $j_{n+1}$, with poles depending only on $j_{n+1}-\sum_{r=1}^{n}j_{r}$. For generic values of $j_{1},\dots,j_{n}$ we may therefore consider the limit $j_{n+1}\to 0$, combined with $E_{n+1}\to 0$, yielding the right side of (LABEL:Fourier(n)).
5.2 Differential equations
We are now going to see that $\Theta^{(n,i)}_{\chi}$ satisfy a holonomic system of differential equations identifying $\Theta^{(n,i)}_{\chi}$ as representations of the eigenfunctions with $z_{n}$ in generic position, and an additional variable $x_{n}$, for $i=0,1$. We will see that the dependence $\Theta^{(n,i)}_{\chi}\big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[2.15277pt] x_{1}&\dots&x_{n}\end{smallmatrix}\big)$ with respect to the variables $x_{1},\dots,x_{n}$ is constrained by certain groups of symmetries extending the symmetry of $\Sigma^{(n,i)}_{\chi}\big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[2.15277pt] x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\big)$ under translations and dilatations. Invariance under the Lie-algebras of the symmetry groups is contained in the following Lemma.
Lemma 4.
The functions $\Theta^{(n,i)}_{\chi}$ defined in (LABEL:Fourier(n)) satisfy the system of equations
$\mathsf{H}_{r}\Theta^{(n,i)}_{\chi}=E_{r}\Theta^{(n,i)}_{\chi},\qquad{\mathcal{D}}^{a}\Theta^{(n,i)}_{\chi}=0,\quad\begin{aligned} &a\in\{0,+\},\\ &i\in\{0,1\},\end{aligned}\qquad\left\{\begin{aligned} &{\mathcal{D}}^{-}\Theta^{(n,0)}_{\chi}=0,\\ &\widehat{\mathcal{D}}^{+}\Theta^{(n,1)}_{\chi}=0,\end{aligned}\right$. (5.47)
where $r\in\{1,\dots,n\}$, $a\in\{-,0,+\}$ and $i\in\{0,1\}$, using the notations
$\mathcal{D}^{0}=\sum_{r=1}^{n}\bigg(x_{r}\frac{\partial}{\partial x_{r}}-j_{r}\bigg)\,,\quad\mathcal{D}^{-}=\sum_{r=1}^{n}\bigg(x_{r}^{2}\frac{\partial}{\partial x_{r}}-2j_{r}x_{r}\bigg),\quad{\mathcal{D}}^{+}=\sum_{r=1}^{n}\frac{\partial}{\partial x_{r}}\,$, (5.48)
$\mathsf{H}_{r}=\sum_{s\neq r}^{n}\frac{{\mathcal{D}}_{r}^{a}{\mathcal{D}}_{s}^{b}\eta_{ab}}{z_{r}-z_{s}},\qquad r=1,\dots,n,\hskip 105.2751pt\widehat{\mathcal{D}}^{+}=\sum_{r=1}^{n}z_{r}\frac{\partial}{\partial x_{r}}.$ (5.49)
Proof.
For the following argument we shall again start by considering $\Sigma^{(n+1)}_{\hat{\chi}}\big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[2.15277pt] x_{1}&\dots&x_{n}\end{smallmatrix};j_{n+1}\big)$, as done at the beginning of this section. $\Sigma^{(n+1)}_{\hat{\chi}}$ satisfies the system of differential equations,
$\mathsf{H}_{r}\Sigma^{(n+1)}_{\chi}=E_{r}\Sigma^{(n+1)}_{\chi},\quad r=1,\dots,{n},\qquad{\mathcal{D}}_{a}\Sigma^{(n+1)}_{\chi}=0,\qquad a\in\{0,+\},$ (5.50)
as follows from Proposition 1. We are going to note that the differential equations stated in Lemma 4 follow from (5.50) in the limit $j_{n+1}\to 0$ and $E_{m+1}\to 0$. The only nontrivial case is the equation ${\mathcal{D}}_{-}\Theta^{(n,0)}_{\chi}=0$.
Let us consider the generating function of differential operators
$\mathsf{T}_{n+1}(z)=\sum_{r=1}^{n}\bigg(\frac{\delta_{r}}{(y-z_{r})^{2}}+\frac{{\mathsf{H}}_{r}}{y-z_{r}}\bigg).$
The Fourier-transformation $\widetilde{\mathsf{T}}_{n+1}(z)$ of $\mathsf{T}_{n+1}(z)$ can be represented in the form (3.27) with $\mathsf{a}(z)$, $\mathsf{b}(z)$ and $\mathsf{c}(z)$ replaced by very similar differential operators, defined by extending the summation over $r$ in the formulae (3.28) to include the value $r=n$. We may note that
$\widetilde{\mathsf{T}}_{n+1}(z)$ $=\frac{1}{z^{2}}(\mathsf{a}_{\infty}^{2}-\mathsf{a}_{\infty}+\mathsf{c}_{\infty}\mathsf{b}_{\infty})+\frac{\rho(\bm{k})}{z^{3}}\,\mathsf{b}_{\infty}+{\mathcal{O}}(z^{-4})$ (5.51)
$=\frac{\delta_{n+1}}{z^{2}}+\frac{E_{n+1}}{z^{3}}+{\mathcal{O}}(z^{-4}),$
using the notations
$\mathsf{a}_{\infty}=-\sum_{r=1}^{n}k_{r}\frac{\partial}{\partial k_{r}},\qquad\mathsf{b}_{\infty}=\sum_{r=1}^{n}\bigg(\frac{\delta_{r}}{k_{r}}-k_{r}\frac{\partial^{2}}{\partial k_{r}^{2}}\bigg),\qquad\mathsf{c}_{\infty}=\sum_{r=1}^{n}k_{r}.$
As the Fourier-transformation defining ${\Theta}^{(n)}_{\chi}$ is supported on $\sum_{r=1}^{n}k_{r}=0$, and as we have $\mathsf{a}_{\infty}\widetilde{\Theta}^{(n)}_{\chi}=j_{n+1}$ by dilatation invariance, we may conclude from (5.51) that $\widetilde{\Theta}^{(n)}_{\chi}=0$ must satisfy $\mathsf{b}_{\infty}\widetilde{\Theta}^{(n)}_{\chi}=0$ in the limit $j_{n+1}\to 0$ and $E_{n+1}\to 0$. This implies the equation ${\mathcal{D}}_{-}\Theta^{(n)}_{\chi}=0$ by Fourier-transformation. ∎
5.3 Relation to inverse SoV-transformation
We will now show that the functions $\Sigma^{(n,i)}_{{\chi}}$ representing the inverse $\mathsf{SoV}$-transformations can be represented as a certain limit of the functions $\Theta^{(n,i)}_{\chi}$ defined in (LABEL:Fourier(n)), for $i=0,1$, respectively. The integrals defining $\Theta^{(n,i)}_{\chi}$ thereby furnish an alternative representation for the inverse $\mathsf{SoV}$-transformations that will be used in the next section.
Lemma 5.
$\lim_{z_{n}\to\infty}|{z_{n}}|^{-2j_{n}}\Theta^{(n,1)}_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[4.09024pt] x_{1}&\dots&x_{n}\end{smallmatrix}\Big)=\Sigma^{(n,1)}_{{\chi}}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[4.09024pt] x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\Big).$ (5.52)
Proof.
We will use the distribution $\delta^{(2)}\big({\textstyle\sum_{r=1}^{n}z_{r}k_{r}}\big)$ in order to perform the integration over $k_{n}$, using the fact that $\sum_{r=1}^{n-1}k_{r}z_{r}$ stays finite in the limit of our interest. It follows that we have
$k_{n}=-z_{n}^{-1}\rho(\bm{k}),\quad\text{and}\quad\hat{\rho}_{1}(\bm{k})=-{z_{n}}\rho(\bm{k})+{\mathcal{O}}(z_{n}^{0}),$ (5.53)
on the support of $\delta^{(2)}\big({\textstyle\sum_{r=1}^{n}z_{r}k_{r}}\big)$, recalling that $\rho(\bm{k})=\sum_{r=1}^{n-1}k_{r}z_{r}$, and $\hat{\rho}_{1}(\bm{k})=\sum_{r=1}^{n}k_{n}z_{n}^{2}$. The relations defining the functions $y_{k}(\bm{k})$ simplify to
$k_{r}=\frac{\hat{\rho}_{1}(\bm{k})}{(z_{r}-z_{n})}\frac{\prod_{k=1}^{n-3}(z_{r}-y_{k}(\bm{k}))}{\prod_{r\neq s}^{n-1}(z_{r}-z_{s})}=\rho(\bm{k})\frac{\prod_{k=1}^{n-3}(z_{r}-y_{k}(\bm{k}))}{\prod_{r\neq s}^{n-1}(z_{r}-z_{s})}+\mathcal{O}(z^{-1}_{n}),$ (5.54)
for $r=1,\dots,n-1$. We can thereby conclude that the functions $y_{k}(\bm{k})$ become independent of $k_{n}$ in the limit $z_{n}\to\infty$. In this way it becomes straightforward to find that the limit of the integral defining $\Theta^{(n,1)}_{\chi}$ on the right side of (5.52) is given by the integral
$\frac{1}{\pi^{n}}\int\prod_{r=1}^{n-1}\frac{d^{2}k_{r}\,e^{2\mathrm{i}\,\mathrm{Im}(k_{r}x_{r})}}{|k_{r}|^{2j_{r}+2}}\delta^{(2)}\big({\textstyle\sum_{r=1}^{n}k_{r}}\big)|\rho(\bm{k})|^{-2j_{n}-2}\prod_{k=1}^{n-3}\chi\big(y_{k}(\bm{k}),\bar{y}_{k}(\bm{k})\big).$ (5.55)
This is the integral defining $\Sigma^{(n,1)}_{{\chi}}\big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[1.07639pt] x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\big)$. ∎
For future use, let us note an interesting consequence of (5.55): We have
$\lim_{z_{n}\to\infty}\Theta^{(n,1)}_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}&-1\\[4.09024pt] x_{1}&\dots&x_{n-1}&x_{n}\end{smallmatrix}\Big)=\Theta^{(n-1,0)}_{{\chi}}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[4.09024pt] x_{1}&\dots&x_{n-1}\end{smallmatrix}\Big).$ (5.56)
This relation will play an important role in our proof of Theorem 3 later.
@@ 5.3 Relation to inverse SoV-transformation
We have
$\lim_{x_{n},z_{n}\to\infty}\Big|\frac{z_{n}}{x_{n}}\Big|^{4j_{n}}\Theta^{(n,0)}_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[4.09024pt] x_{1}&\dots&x_{n}\end{smallmatrix}\Big)=\chi_{n}\,\frac{\pi\,\Gamma(-2j_{n})}{\Gamma(2j_{n}+1)}\Sigma^{(n,0)}_{{\chi}}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[4.09024pt] x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\Big),$ (5.57)
with $\chi_{n}$ being a short notation for $C_{n,-}(\bm{j})$ defined in (3.33).
Proof.
In order to verify (5.57) let us note that the leading contributions to the limit $x_{n}\to\infty$ are coming from the region where $k_{n}\to 0$. It follows from (5.46) that there must exist an index $k$ such that $y_{k}\to z_{n}$. The permutation symmetry of the integrand allows us to assume without loss of generality that $y_{n-2}\to z_{n}$. We may next note that
$k_{r}=\rho\frac{(z_{r}-y_{n-2})}{(z_{r}-z_{n})}\frac{\prod_{k=1}^{n-3}(z_{r}-y_{j})}{\prod_{r\neq s}^{n-1}(z_{r}-z_{s})}=\rho\frac{\prod_{k=1}^{n-3}(z_{r}-y_{k})}{\prod_{r\neq s}^{n-1}(z_{r}-z_{s})}+\mathcal{O}(x^{-1}_{n}).$ (5.58)
It follows that the functions $y_{k}(\bm{k})$, $k=1,\dots,n-3$, and $\hat{\rho}_{0}(\bm{k})$ become independent of $k_{n}$ in the limit of our interest.
In the limit $y_{n-2}\to z_{n}$ we may furthermore use
$k_{n}=\rho{(z_{n}-y_{n-2})}\frac{\prod_{k=1}^{n-3}(z_{n}-y_{j})}{\prod_{s=1}^{n-1}(z_{n}-z_{s})},$ (5.59)
together with the asymptotic behaviour (3.33). As $k_{n}$ is proportional to $y_{n-2}-z_{n}$, we see that the dependence of the integrand defining the right side of (5.57) with respect to $k_{n}$ simplifies significantly. We may factorise the Fourier-transformation over $k_{1},\dots,k_{n}$ into an integration over $k_{n}$, followed by a Fourier-transformation over $k_{1},\dots,k_{n-1}$. The integration over $k_{n}$ can be carried out explicitly using the identity (see e.g. equation (7) of Appendix B1.7 in [GeS])
$\frac{2j+1}{\pi}\int_{\mathbb{C}}d^{2}x^{\prime}\;|x-x^{\prime}|^{-4j-4}e^{kx^{\prime}-\bar{k}\bar{x}^{\prime}}=e^{kx-\bar{k}\bar{x}}|k|^{2(2j+1)}\frac{\Gamma(-2j-1)}{\Gamma(2j+1)}\,.$ (5.60)
One of the resulting terms is a distribution supported at $x_{n}=0$ which can not contribute to the limit $x_{n}\to\infty$. The other term has an asymptotic behaviour for $z_{n}\to\infty$ of the following form
$\frac{\chi_{n}}{\pi^{n}}\left|\frac{x_{n}}{z_{n}}\right|^{4j_{n}}\frac{\pi\,\Gamma(-2j_{n})}{\Gamma(2j_{n}+1)}\int\prod_{r=1}^{n-1}\frac{d^{2}k_{r}\,e^{2\mathrm{i}\,\mathrm{Im}(k_{r}x_{r})}}{|k_{r}|^{2j_{r}+2}}\delta^{(2)}\big({\textstyle\sum_{r=1}^{n-1}}k_{r}\big)|\rho(\bm{k})|^{2(j_{n}+1)}\prod_{k=1}^{n-3}\chi\big(y_{k}(\bm{k}),\bar{y}_{k}(\bm{k})\big),$
which is proportional to the left side of (5.57). ∎
By solving the differential equations satisfied by $\Theta^{(n,i)}_{\chi}$ with the boundary condition provided by (5.57) one may express the functions $\Theta^{(n)}_{\chi}$ defined in (LABEL:Fourier(n)) in terms of $\Sigma^{(n)}_{\chi}\big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[2.15277pt] x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\big)$.
6 Hecke modifications of the inverse SoV transformation
We shall here prove the crucial Theorems 3 and 4. We will consider the inverse $\mathsf{SoV}$-transformations associated to Riemann spheres $C_{0,n+1}$ obtained from $C_{0,n}$ by adding another marked point at $z=z_{0}$. When the parameter $j_{0}$ associated to $z_{0}$ is chosen as $j_{0}=-1$, we will find a relation with the Hecke modification $\mathbb{H}_{z_{0},x_{0}}$, and with the Hecke operator $\mathsf{H}_{z_{0}}$.
6.1 Differential equations
It will be crucial to observe that the system of differential equations satisfied by $\Theta^{(n+1,1)}_{\chi}$ simplifies somewhat when $j_{0}=-1$, and $E_{0}=0$.
Proposition 4.
When $j_{0}=-1$ and $E_{0}=0$, the distribution $\Theta^{(n+1,1)}_{\chi}$ satisfies the differential equations ${\mathcal{D}}_{a}\Theta^{(n+1,1)}_{\chi}=0$, $a\in\{-,0,+\}$, and furthermore
@@ 6.1 Differential equations
with $\widetilde{\Theta}^{(n+1,1)}_{\chi}$ being the Fourier-transformation of ${\Theta}^{(n+1,1)}_{\chi}$, with $\widetilde{\mathsf{T}}(z)$ defined in (3.27), and
@@
It follows that $\widetilde{\mathcal{N}}\widetilde{\Theta}^{(n+1,1)}_{\chi}=0$, implying (6.61) by inverse Fourier-transformation. ∎
@@
6.2 Proof of Theorems 3 and 4

### v2

@@ 5 Hecke modifications of the inverse SoV transformation
$\Sigma^{(n+1,1)}_{\chi}\Big(\begin{smallmatrix}j_{0}&j_{1}&\dots&j_{n-1}\\[3.22916pt] x_{0}&x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\Big)=\int\prod_{r=0}^{n-1}\frac{d^{2}k_{r}\,e^{2\mathrm{i}\,\mathrm{Im}(k_{r}x_{r})}}{\pi\,|k_{r}|^{2j_{r}+2}}$ $\pi\delta^{(2)}\big({\textstyle\sum_{r=0}^{n-1}k_{r}}\big)$ (5.47)
$\times|\hat{\rho}(\bm{k})|^{-2j_{n}}\prod_{k=0}^{n-3}\chi\big(y_{k}(\bm{k}),\bar{y}_{k}(\bm{k})\big),$
with $y_{0}(\bm{k}),\dots,y_{n-3}(\bm{k})$ and $\hat{\rho}(\bm{k})$, $\bm{k}=(k_{0},\dots,k_{n-1})$, defined by
$k_{r}=\hat{\rho}_{i}(\bm{k})\frac{\prod_{k=0}^{n-3}(z_{r}-y_{k}(\bm{k}))}{\prod_{s\neq r}(z_{r}-z_{s})},\qquad r=0,\dots,n-1,\qquad\hat{\rho}(\bm{k})=\sum_{r=0}^{n-1}z_{r}k_{r}.$ (5.48)
Theorem 3 is equivalent to the following integral identity:
$\Sigma^{(n+1,1)}_{\chi}\Big(\begin{smallmatrix}-1&j_{1}&\dots&j_{n-1}\\[3.22916pt] x_{0}&x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\Big)=\,\chi_{n}\,\prod_{r=1}^{n-1}\frac{|x_{r}-x_{0}|^{4j_{r}}}{|z_{r}-z_{0}|^{2j_{r}}}\,\Sigma^{(n,0)}_{{\chi}}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[1.07639pt] -\frac{z_{1}-z_{0}}{x_{1}-x_{0}}&\dots&-\frac{z_{n-1}-z_{0}}{x_{n-1}-x_{0}}\end{smallmatrix};j_{n}\Big),$ (5.49)
with function $\Sigma^{(n,0)}_{{\chi}}=\big(\mathsf{S}_{\bm{j},0}^{(n)}\big)^{-1}\phi_{\chi}$ on the right side of (5.49) being represented by an integral obtained from (5.47) by restricting the ranges for $r$ and $k$ to $\{1,\dots,n-1\}$ and $\{1,\dots,n-3\}$, respectively, and replacing $j_{n}$ by $j_{n}-1$. In order to prove (5.49) we will first derive a differential equations for the function $\Sigma^{(n+1,1)}$ on the left side of (5.49), then note that the right side of (5.49) satisfies the same equations, and finally establish the equality of the limits $z_{0}\to 0$, followed by $x_{0}\to 0$.
5.1 Differential equation
It will be crucial to observe that one of the eigenvalue equations satisfied by $\Theta^{(n+1,1)}_{\chi}$ simplifies to a first order differential equation when $j_{0}=-1$, and $E_{0}=0$.
@@ 5.1 Differential equation
When $j_{0}=-1$ and $E_{0}=0$, the distribution $\Theta^{(n+1,1)}_{\chi}$ satisfies
@@
with $\widetilde{\Theta}^{(n+1,1)}_{\chi}$ being the Fourier transformation of ${\Theta}^{(n+1,1)}_{\chi}$, with $\widetilde{\mathsf{T}}(z)$ defined in (3.28), and
@@
It follows that $\widetilde{\mathcal{N}}\widetilde{\Theta}^{(n+1,1)}_{\chi}=0$, implying (5.50) by inverse Fourier transformation. ∎
@@
It follows from the differential equations that there must exist a function $\Psi_{\chi}$ satisfying (2.1) such that
$\Sigma^{(n+1,1)}_{\chi}\Big(\begin{smallmatrix}-1&j_{1}&\dots&j_{n-1}\\[3.22916pt] x_{0}&x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\Big)=\prod_{r=1}^{n-1}\frac{|x_{r}-x_{0}|^{4j_{r}}}{|z_{r}-z_{0}|^{2j_{r}}}\,\Psi_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[1.07639pt] -\frac{z_{1}-z_{0}}{x_{1}-x_{0}}&\dots&-\frac{z_{n-1}-z_{0}}{x_{n-1}-x_{0}}\end{smallmatrix};j_{n}\Big).$ (5.51)
In order to prove (5.49), it remains to compare the limit $z_{0}\to\infty$ followed by $x_{0}\to\infty$ of both sides of equation (5.49).
5.2 Comparison of limits
Note, on the one hand, that we have
$\lim_{z_{0}\to\infty}$ $|z_{0}|^{2j_{n}}\prod_{r=1}^{n-1}{|z_{r}-z_{0}|^{-2j_{r}}}\Psi_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[1.07639pt] -\frac{z_{1}-z_{0}}{x_{1}-x_{0}}&\dots&-\frac{z_{n-1}-z_{0}}{x_{n-1}-x_{0}}\end{smallmatrix};j_{n}\Big)$
$=\lim_{z_{0}\to\infty}|z_{0}|^{-2J_{0}}\Psi_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[1.07639pt] \frac{z_{0}}{x_{1}-x_{0}}&\dots&\frac{z_{0}}{x_{n-1}-x_{0}}\end{smallmatrix};j_{n}\Big)=\Psi_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[1.07639pt] \frac{1}{x_{1}-x_{0}}&\dots&\frac{1}{x_{n-1}-x_{0}}\end{smallmatrix};j_{n}\Big),$
and using $\frac{x_{0}^{2}}{x_{0}-x_{r}}=x_{0}+x_{r}+{\mathcal{O}}(x_{0}^{-1})$ furthermore
$\lim_{x_{0}\to\infty}|x_{0}|^{-4j_{n}}\prod_{r=1}^{n-1}{|x_{r}-x_{0}|^{4j_{r}}}\Psi_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[1.07639pt] \frac{1}{x_{1}-x_{0}}&\dots&\frac{1}{x_{n-1}-x_{0}}\end{smallmatrix};j_{n}\Big)$
$=\lim_{x_{0}\to\infty}\Psi_{\chi}\bigg(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[2.15277pt] \frac{x_{0}^{2}}{x_{0}-x_{1}}&\dots&\frac{x_{0}^{2}}{x_{0}-x_{n-1}}\end{smallmatrix};j_{n}\bigg)=\lim_{x_{0}\to\infty}\Psi_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[3.22916pt] x_{0}+x_{1}&\dots&x_{0}+x_{n-1}\end{smallmatrix};j_{n}\Big)=\Psi_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[3.22916pt] x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\Big).$
In order to prove Theorem 3 one need to compare this with the limits of the left side of (5.49).
Proposition 4.
We have
$\lim_{z_{0}\to\infty}|z_{0}|^{2j_{n}}\Sigma^{(n+1,1)}_{\chi}\Big(\begin{smallmatrix}j_{0}&j_{1}&\dots&j_{n-1}\\[3.22916pt] x_{0}&x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\Big)=\Theta^{(n,0)}_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}&j_{n}\\[3.22916pt] x_{1}&\dots&x_{n-1}&x_{0}\end{smallmatrix}\Big),$ (5.52)
$\lim_{x_{n}\to\infty}|{x_{n}}|^{-4j_{n}}\Theta^{(n,0)}_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[4.09024pt] x_{1}&\dots&x_{n}\end{smallmatrix}\Big)=\chi_{n}\,\Sigma^{(n,0)}_{{\chi}}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[4.09024pt] x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\Big),$ (5.53)
with $\chi_{n}$ being defined in (4.44), and
$\Theta^{(n,0)}_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}&j_{n}\\[3.22916pt] x_{1}&\dots&x_{m-1}&x_{n}\end{smallmatrix}\Big)=\int\prod_{r=1}^{n}\frac{d^{2}k_{r}\,e^{2\mathrm{i}\,\mathrm{Im}(k_{r}x_{r})}}{\pi\,|k_{r}|^{2j_{r}+2}}$ $\;\pi\delta^{(2)}\big({\textstyle\sum_{r=1}^{n}k_{r}}\big)$ (5.54)
$\times\big|\theta(\bm{k})\big|^{2}\,\prod_{k=0}^{n-3}\chi\big(y_{k}(\bm{k}),\bar{y}_{k}(\bm{k})\big),$
with $y_{0}(\bm{k}),\dots,y_{n-3}(\bm{k})$ and $\theta(\bm{k})$, $\bm{k}=(k_{1},\dots,k_{n})$, defined by
$k_{r}=\theta(\bm{k})\frac{\prod_{k=0}^{n-3}(z_{r}-y_{k}(\bm{k}))}{\prod_{s\neq r}(z_{r}-z_{s})},\qquad r=1,\dots,n,\qquad\theta(\bm{k})=\sum_{r=1}^{n-1}k_{r}.$ (5.55)